OpenStudy (anonymous):

ernest has a job helping a jewler. he had the assignment of counting faces, vertices and edges on the emeralds. on the first emerald enrest hcounted 6 faces and 6 vertices. he quickly realized he didnt have to count the edges. how many edges were there?

OpenStudy (anonymous):

This can be easily calculated using Eular's Formula V + F - E =2

OpenStudy (anonymous):

Here V = number of vertices F = number of faces and E = number of edges

OpenStudy (anonymous):

In your case, F = 6 V = 6 Substituting them in Eular's formula, we get 6 + 6 - E = 2 12 - E = 2 12 - 2 = E 10 = E Hope this is clear??? :D

OpenStudy (anonymous):

yes, thanks sooo much!

OpenStudy (anonymous):

Got a medal for me ???? :D

OpenStudy (anonymous):

new to this whole thing lol how might i do that?

OpenStudy (anonymous):

Just look on the right side You'll see a icon "Give Medal" Click on that to giv someone a medal if u feel they deserve it

OpenStudy (anonymous):

i dont see it : (

OpenStudy (anonymous):

im sorry. u have a medal in my book! lol i keep trying to figure it out

OpenStudy (anonymous):

Just above my reply, there is a bar showing my name Look at its extreme right side

OpenStudy (anonymous):

its says 25 medals but wont let me click is. i became a fan. hope that helps. : )

OpenStudy (anonymous):

It really doesn't matter. Just shows appreciation for those who help you........